# Backdoor Attack with Sparse and Invisible Trigger

### v2

@@ IV-E4 SIBA with Limited Training Data
IV-E5 SIBA with Asymmetric Triggers
Figure 15: Results of SIBA with amplified triggers on CIFAR-10.
@@ V Conclusion
In this paper, we proposed a novel backdoor attack, $i.e.$, sparse and invisible backdoor attack (SIBA), to achieve attack effectiveness and attack stealthiness simultaneously. Our SIBA method only needs to modify a few pixels of the original images to generate poisoned samples and is human-imperceivable due to the low modification magnitude. To achieve it, we formulated the trigger generation as a bi-level optimization problem with sparsity and invisibility constraints and proposed an effective method to solve it. We conducted extensive experiments on benchmark datasets, verifying the effectiveness, the resistance to potential defenses, and the flexibility under different settings of our attack. We hope our method can provide a new angle and deeper understanding of backdoor mechanisms, to facilitate the design of more secure and robust DNNs.
@@
[…]

### v3

@@ IV-E4 SIBA with Limited Training Data
IV-E5 SIBA with Asymmetric Triggers
IV-E6 Effectiveness of SIBA under Clean-label Settings
To evaluate the effectiveness of SIBA under clean-label settings, we conduct experiments on the CIFAR-10 dataset. It is known that clean-label attacks are more challenging than poisoned-label ones and we thus set $k=200,\epsilon=16/255$. The results are summarized in Table VI, from which we could find that our SIBA achieves $>90\%$ ASR with $5\%$ poisoning rate and outperforms the other baselines with a notable margin. However, we have to acknowledge that when the sparsity is further reduced ($e.g.$, $k=100$), clean-label settings may not achieve high ASRs like poisoned-label ones. We will explore more advanced optimization techniques to push the limit of sparsity under clean-label settings in our future work.
Figure 15: Results of SIBA with amplified triggers on CIFAR-10.
TABLE VI: Comparison with the baseline attacks under clean-label settings on the CIFAR-10 dataset.
[…]
Random BA (%) 94.51 94.42 94.56 94.10 93.87
ASR (%) 1.51 4.91 45.58 68.89 79.81
Sparse BA (%) 94.48 94.80 94.26 94.34 94.06
ASR (%) 24.76 46.49 59.87 61.22 73.47
SIBA BA (%) 94.44 94.22 94.45 94.40 94.17
ASR (%) 62.35 75.57 76.89 86.30 91.73
@@ V Conclusion
In this paper, we proposed a novel backdoor attack, $i.e.$, sparse and invisible backdoor attack (SIBA), to achieve attack effectiveness and attack stealthiness simultaneously. Our SIBA method only needs to modify a few pixels of the original images to generate poisoned samples and is human-imperceptible due to the low modification magnitude. To achieve it, we formulated the trigger generation as a bi-level optimization problem with sparsity and invisibility constraints and proposed an effective method to solve it. We conducted extensive experiments on benchmark datasets, verifying the effectiveness, the resistance to potential defenses, and the flexibility under different settings of our attack. We hope our method can provide a new angle and deeper understanding of backdoor mechanisms, to facilitate the design of more secure and robust DNNs.
Acknowledgments
The work is supported in part by the National Natural Science Foundation of China (U20A20178, 62171248, U20B2049, and U21B2018), Shenzhen Science and Technology Program (JCYJ20220818101012025), and the PCNL KEY project (PCL2023AS6-1). This work was mostly done when Yiming Li was a Research Professor at Zhejiang University. He is currently a Research Fellow at Nanyang Technological University.
@@ Acknowledgments
${\bm{t}}_{i+1,j}=\left\{\begin{aligned} &{\bm{v}}_{i,j}&\text{if}\ j\in C^{\prime}\\ &0&\text{if}\ j\notin C^{\prime}\end{aligned},\right.$ (7)